OpenStudy (anonymous):

Medal and fan ASAP |2x-1|-4=3

OpenStudy (anonymous):

help me understand?

pooja195 (pooja195):

add the 4 to both sides

pooja195 (pooja195):

That way it will be easier to solve this and we can isolate the statement from any numbers out of the bars

OpenStudy (anonymous):

|2x-1|=7

OpenStudy (anonymous):

then?

pooja195 (pooja195):

now set up 2 equations... do you know what the equations would look like/

OpenStudy (anonymous):

no?

pooja195 (pooja195):

We know either \[\huge ~\rm~2x−1=7 ~~~~~~or ~~~~~~2x−1=−7\]

OpenStudy (anonymous):

ok.. next?

pooja195 (pooja195):

solve and you will end up with TWO solutuions

OpenStudy (anonymous):

how? Its just absolute value? the answer is always positive...

OpenStudy (anonymous):

i just need step by step how to

OpenStudy (anonymous):

next???

pooja195 (pooja195):

We need to solve the equations now. We set up to equations to show that \[\huge~\rm~|2x-1|\] either equations positive 7 or negative 7

OpenStudy (anonymous):

ok.. next?

pooja195 (pooja195):

We know either \[\huge ~\rm~2x−1=7 ~~~~~~or ~~~~~~2x−1=−7\]

OpenStudy (anonymous):

ok.. next

pooja195 (pooja195):

solve

OpenStudy (anonymous):

x=-3 or x=4

OpenStudy (anonymous):

|2x-1|+4=3

OpenStudy (anonymous):

2x-1=-1 ??

OpenStudy (anonymous):

not solvable!!?

pooja195 (pooja195):

\(\color{blue}{\text{Originally Posted by}}\) @Adrianna765 x=-3 or x=4 \(\color{blue}{\text{End of Quote}}\) this is right

OpenStudy (anonymous):

ok.. thanks!!!